# Quantum metrology with a continuous-variable system

### v1

@@ III.4 Classical Fisher information
$F_{\mathrm{G}}(\theta)=\frac{1}{\sigma^{2}}\left(\left(\frac{\partial\mu}{\partial\theta}\right)^{2}+\frac{1}{2\sigma^{2}}\left(\frac{\partial\sigma^{2}}{\partial\theta}\right)^{2}\right).$ (49)
@@
$F[\hat{\rho}(\theta),\hat{G},\bm{w}^{T}\mathbf{\hat{r}}]=\frac{\left(\bm{w}^{T}\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{2}}{\bm{w}^{T}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\bm{w}}+\frac{\left(\bm{w}^{T}\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}\bm{w}\right)^{2}}{2(\bm{w}^{T}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\bm{w})^{2}}.$ (50)
@@ IV Displacement sensing
The first class of perturbations we are going to consider are the one resulting in a phase-space translation of the system’s state. These can be expressed by the displacement operator Eq. (4), that we rewrite here as
@@ IV.1.3 Arbitrary Gaussian states
$F_{Q}^{\max}[\hat{\rho}]$ $=2\mathcal{P}(\hat{\rho})^{2}((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}+\sqrt{4\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}+((\Delta\hat{x})^{2}_{\hat{\rho}}-(\Delta\hat{p})^{2}_{\hat{\rho}})^{2}})\;,$ (68)
$F_{Q}^{\min}[\hat{\rho}]$ $=2\mathcal{P}(\hat{\rho})^{2}((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}-\sqrt{4\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}+((\Delta\hat{x})^{2}_{\hat{\rho}}-(\Delta\hat{p})^{2}_{\hat{\rho}})^{2}})\;,$ (69)
@@
$F_{Q}^{\text{avg}}[\hat{\rho}]$ $=2\mathcal{P}(\hat{\rho})^{2}((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}})\;.$ (70)
@@ IV.1.4 Performance of different quantum states
Note that, taking coherent states as a classical benchmark [68], any value of $F_{Q}[\hat{\rho},\hat{r}(\phi)]>2$ would indicate a quantum enhancement in sensing displacements.
@@
Coherent $|\alpha\rangle$ $2$
@@
Coherent $|\alpha\rangle$ $2$ $2$ $2$
@@
Fock superposition $(|m\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$ $n>m$$\begin{cases}4n\qquad\qquad\qquad&n=m+1\\ 2((2n-1)+\sqrt{n(n-1)})&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$$\begin{cases}2n\qquad\qquad\qquad&n=m+1\\ 2((2n-1)-\sqrt{n(n-1)})&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$$\begin{cases}3n\qquad\qquad&n=m+1\\ 2(2n-1)\qquad\qquad&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$
Cat $\left|\alpha\right|\gtrsim 2$ $2(1+4\overline{n})$ $2$ $2(1+2\overline{n})$
@@
Table 2: Maximum, minimum and average QFI for a displacement along the direction specified by $\phi$. For a definition of these quantities, see Eqs. (53,54,55).
@@ IV.2.1 Homodyne measurements
and analogously for $\max_{\varepsilon}(\chi^{-2})^{\min}_{\mathrm{hom,fix}}$ and $\max_{\varepsilon}(\chi^{-2})^{\mathrm{avg}}_{\mathrm{hom,fix}}$ with prefactors 00 and $1/2$, respectively. Since this maximum over $\varepsilon$ is achieved by choosing $\bm{w}$ along the direction of minimal phase-space variance, this means that, if $\phi$ is unknown, the above figures of merit are maximized by the measurement with the smallest variance. Similarly, we could easily obtain the minimum or average over $\varepsilon$. The evaluation of $(\chi^{-2})^{\max}_{\mathrm{hom,fix}}[\hat{\rho},\hat{r}(\varepsilon)]$ and $(\chi^{-2})^{\max}_{\mathrm{hom,max}}[\hat{\rho}]$ for a selection of relevant families of states can be found in Table 4.
@@
Coherent $|\alpha\rangle$ $2$ $2$
@@
Fock superposition $(|0\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$$\begin{cases}\frac{1}{1-\frac{1}{2}\sin^{2}(\gamma+\varepsilon)}\qquad&n=1\\ \frac{1}{\frac{3}{2}-\frac{1}{\sqrt{2}}\cos(\gamma+2\varepsilon)}&n=2\\ \frac{2}{1+n}&\text{otherwise}\\ \end{cases}$$\begin{cases}2\qquad&n=1\\ 2\left(3+\sqrt{2}\right)/7&n=2\\ \frac{2}{1+n}&\text{otherwise}\\ \end{cases}$
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$, $\alpha\in\mathbb{C}$, $\left|\alpha\right|\gtrsim 2$ $\dfrac{2}{1+4\overline{n}^{2}\cos^{2}\varepsilon}$ $2$
@@
Table 4: Displacement sensitivity attainable by linear quadrature measurements. Sensitivities computed according to Eqs. (76,79), corresponding to cases where the measurement direction $\varepsilon$ is chosen before the generator is known.
@@
$(\chi^{-2})^{\mathrm{avg}}_{\mathrm{hom}}[\hat{\rho}]$ $=\frac{1}{2\pi}\int_{0}^{2\pi}\text{d}\phi\,\chi^{-2}_{\mathrm{hom}}[\hat{\rho},\hat{r}(\phi)]=\frac{\frac{1}{2}((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}})}{(\Delta\hat{x})^{2}_{\hat{\rho}}(\Delta\hat{p})^{2}_{\hat{\rho}}-\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}}\;,$ (85)
@@
The most relevant quantity above, $(\chi^{-2})^{\max}_{\mathrm{hom}}[\hat{\rho}]$, expresses the maximal sensitivity that can be obtained from a quantum state $\hat{\rho}$ under displacements with homodyne measurements. The maximization over $\phi$ describes the optimal orientation of the displacement with respect to the fluctuations of the state. This is achieved by choosing $\phi$ such that $\bm{u}$ coincides with the maximal eigenvector of $\Gamma[\hat{\rho},\mathbf{\hat{r}}]$. In the case of a squeezed state, for instance, this means that the displacement is realized in the direction of the squeezed quadrature and, therefore, the generator must coincide with the anti-squeezed quadrature. The optimization over $\varepsilon$, Eq. (81), assures that the homodyne measurement is realized with an optimal orientation. It is interesting to note that if the displacement direction $\bm{u}$ coincides with an eigenvector of $\Gamma[\hat{\rho},\mathbf{\hat{r}}]$, as is the case for the maximal sensitivity Eq. (83), the optimal homodyne direction $\bm{w}$ is orthogonal to the displacement direction $\bm{u}$. To see this, note that from Eq. (47) follows that the optimal direction $\bm{w}=\zeta\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega\bm{u}_{i}=\zeta^{\prime}\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\bm{u}_{i}=\zeta^{\prime\prime}\Omega\bm{u}_{i}$ is indeed orthogonal to $\bm{u}_{i}$, where $\Gamma[\hat{\rho},\mathbf{\hat{r}}]\bm{u}_{i}=\lambda_{i}(\Gamma[\hat{\rho},\mathbf{\hat{r}}])\bm{u}_{i}$, $\zeta^{\prime\prime}=\lambda_{i}(\Gamma[\hat{\rho},\mathbf{\hat{r}}])\zeta^{\prime}$, $\zeta^{\prime}=\zeta/(\det\Gamma[\hat{\rho},\mathbf{\hat{r}}])$ and $\zeta$ are irrelevant real constants, and we used that $\Gamma[\hat{\rho},\mathbf{\hat{r}}]$ is a $2\times 2$ matrix.
@@
Coherent $|\alpha\rangle$ $2$ $2$ $2$
@@
Fock superposition $(|0\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$$\begin{cases}2\qquad&n=1\\ 2(3+\sqrt{2})/7&n=2\\ 2/(1+n)&\text{otherwise}\end{cases}$$\begin{cases}\dfrac{2}{(3+\sqrt{2})}\qquad&n=2\\ \dfrac{2}{(1+n)}&\text{otherwise}\end{cases}$$\begin{cases}3/2\qquad&n=1\\ 6/7&n=2\\ 2/(1+n)&\text{otherwise}\end{cases}$
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$ $\alpha\in\mathbb{C}$, $\left|\alpha\right|\gtrsim 2$ $2$ $\dfrac{2}{1+4\overline{n}}$ $\dfrac{2+4\overline{n}}{1+4\overline{n}}$
Compass $\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$ $\dfrac{2}{1+2\overline{n}}$ $\dfrac{2}{1+2\overline{n}}$ $\dfrac{2}{1+2\overline{n}}$
Table 5: Displacement sensitivity attainable by linear quadrature measurements. Sensitivities computed according to Eqs. (83,84,85), corresponding to cases where the measurement direction $\varepsilon$ is chosen after the generator is known.
@@ IV.2.2 Photon number measurements
where we used that $[\hat{n},\hat{x}]=-i\hat{p}$ and $[\hat{n},\hat{p}]=i\hat{x}$.
@@
[…]
@@ IV.2.3 Measurement of higher-order moments
[…]
@@
$\hat{M}_{\mathrm{opt}}=\cos\phi\left[-(2n+1)\hat{x}+\hat{x}^{3}+\frac{\hat{x}\hat{p}^{2}+\hat{p}\hat{x}\hat{p}+\hat{p}^{2}\hat{x}}{3}\right]-\sin\phi\left[-(2n+1)\hat{p}+\hat{p}^{3}+\frac{\hat{x}^{2}\hat{p}+\hat{x}\hat{p}\hat{x}+\hat{p}\hat{x}^{2}}{3}\right]\;.$ (92)
@@ IV.2.5 Parity operators and Wigner function measurements
according to Eq. (43), the moment-based sensitivity from its average value coincides with the classical Fisher information of the full probability for obtaining the results $1$ or $-1$ respectively,
@@
We note that there are points in phase space where Eq. (100) is indeterminate. The reason for this is because the Wigner function is a bounded function, and therefore both numerator and denominator of Eq. (100) are zero at its extremal points. In other words, these indeterminate values happen at points in phase space where a parity measurement gives ±$\pm 1$, such as for Fock states. However, we note that this indeterminate sensitivity is a mathematical artifact, that cannot be observed in experiments. In practice there will always be some finite amount of noise that prevents from observing $(\Delta\hat{M}(\beta))^{2}_{\hat{\rho}}=0$, and thus from having a zero denominator in Eq. (100). To describe this effect we add to the denominator of Eq. (100) a small positive number $\varepsilon$ representing measurement noise, namely we write
@@
Figure 3 shows the displacement sensitivity for parity measurements of Fock states, while Figures 4, 5 discuss superpositions of two Fock states. Figure 6 is for Gaussian states, while Figure 7 is for cats and compass states.
Figure 3: Displacement sensitivity of Fock states for parity measurements. Left panel: slices for $p=0$ of the Wigner functions for the first four Fock states. Right panel: sensitivity calculated from Eq. (100) (solid line), compared to the QFI bound $24n+$ 2 (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line).
Figure 4: Displacement sensitivity of superpositions of two Fock states for parity measurements. Top row: Wigner functions of different Fock state superpositions. Middle row: sensitivity to displacements obtained through Eq. (100). The displacement direction is optimized in order to maximise the sensitivity, and is indicated by coloured lines. The black dot indicate the point of maximum sensitivity (note that states with discrete rotation symmetry have several points with equal maximum sensitivity). Bottom row: sensitivity along the slices indicated by the lines in the two top rows calculated from Eq. (100) (solid line), compared to the QFI bound $24n+$ 2 (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line). Note that these two plots are not symmetric around zero, and that the QFI bounds are saturated for $n$ even or odd respectively, see Fig. 5.
Figure 5: Displacement sensitivity of superpositions of two Fock states for parity measurements. Sensitivity calculated from Eq. (100) (solid line), compared to the QFI bound $24n+$ 2 (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line). Note that the QFI bounds are saturated for $n$ even or odd respectively.

### v2

@@ III.4 Classical Fisher information
$F_{\mathrm{G}}(\theta)=\frac{1}{\sigma^{2}}\left[\left(\frac{\partial\mu}{\partial\theta}\right)^{2}+\frac{1}{2\sigma^{2}}\left(\frac{\partial\sigma^{2}}{\partial\theta}\right)^{2}\right].$ (53)
@@
$F[\hat{\rho}(\theta),\bm{w}^{T}\mathbf{\hat{r}}]=\frac{\left(\bm{w}^{T}\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{2}}{\bm{w}^{T}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\bm{w}}+\frac{\left(\bm{w}^{T}\frac{\partial\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]}{\partial\theta}\bm{w}\right)^{2}}{2(\bm{w}^{T}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]\bm{w})^{2}}.$ (54)
@@ IV Displacement sensing
Figure 4: Phase-space illustration of a displacement. The metrological task considered in this section is to estimate the displacement amplitude $\theta$.
The first class of perturbations we are going to consider are the one resulting in a phase-space translation of the system’s state, see Fig. 4. These can be expressed by the displacement operator Eq. (4), that we rewrite here as
@@ IV.1.3 Arbitrary Gaussian states
$F_{Q}^{\max}[\hat{\rho}$] $=2\mathcal{P}(\hat{\rho})^{2}\left((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}+\sqrt{4\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}+((\Delta\hat{x})^{2}_{\hat{\rho}}-(\Delta\hat{p})^{2}_{\hat{\rho}})^{2}}\right)\;,$ (72)
$F_{Q}^{\min}[\hat{\rho}$] $=2\mathcal{P}(\hat{\rho})^{2}\left((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}-\sqrt{4\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}+((\Delta\hat{x})^{2}_{\hat{\rho}}-(\Delta\hat{p})^{2}_{\hat{\rho}})^{2}}\right)\;,$ (73)
@@
$F_{Q}^{\text{avg}}[\hat{\rho}$] $=2\mathcal{P}(\hat{\rho})^{2}\left((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}\right)\;.$ (74)
@@ IV.1.4 Performance of different quantum states
The standard quantum limit (SQL) is set by the sensitivity of an optimally used coherent state. For displacement sensing, this yields a classical limit of $F_{Q}[|\alpha\rangle,\hat{r}(\phi)]=2$, independent of $\mathrm{bar}{n}$. Any value $F_{Q}[\hat{\rho},\hat{r}(\phi)]>2$ thus indicates a quantum enhancement in sensing displacements. therefore signifies quantum-enhanced displacement sensitivity. It also certifies the nonclassicality of the state $\hat{\rho}$ as such performance cannot be reproduced by a convex mixture of coherent states, i.e. by a classical Glauber–Sudarshan $P$-representation [79].
@@
Coherent $|\alpha\rangle$ 22
@@
Coherent $|\alpha\rangle$ 22 22 22
@@
Fock superposition
$(|m\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$
$n>m$
$\begin{cases}4n\qquad\qquad\qquad&n=m+1\\ 2((2n-1)+\sqrt{n(n-1)})&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$ $\begin{cases}2n\qquad\qquad\qquad&n=m+1\\ 2((2n-1)-\sqrt{n(n-1)})&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$ $\begin{cases}3n\qquad\qquad&n=m+1\\ 2(2n-1)\qquad\qquad&n=m+2\\ 2(1+n+m)&\text{otherwise}\end{cases}$
Cat $\left|\alpha\right|\gtrsim 2$ $(1+4\overline{n}) 22$ $2(1+2\overline{n})$
@@
Table 2: Maximum, minimum and average QFI for a displacement along the direction specified by $\phi$. For a definition of these quantities, see Eqs. (57,58,59).
IV.1.5 Noisy displacement sensing
If decoherence occurs during interrogation, the QFI, and thus the displacement sensitivity, is typically reduced. The exact dependence on noise and initial state is generally nontrivial, but for Gaussian states under diffusive Gaussian noise, an analytical expression is available. This is particularly relevant, because it describes the experimentally common scenario of coupling to a thermal bath with average photon number $n_{b}$ and decay rate $\kappa$ (see Appendix F.2.3). For an initially pure Gaussian state, we obtain the maximal QFI as a function of the evolution time $t$:
$F_{Q}^{\text{max}}[\hat{\rho},t]=\dfrac{2\left(1+2\overline{n}+2\sqrt{\overline{n}(1+\overline{n})}\right)}{\left((1+2n_{b})e^{\kappa t/2}-2n_{b}e^{-\kappa t/2}\right)^{2}+4(1-e^{-\kappa t})(1+2n_{b})\overline{n}}\;.$ (75)
This expression equals the optimal noiseless QFI Eq. (IV.1.2) multiplied by a reduction factor not exceeding unity for $t,\kappa,n_{b}\geq 0$. Because this factor depends linearly on $\mathrm{bar}{n}$, it suppresses the ideal linear scaling with $\mathrm{bar}{n}$ and drives the QFI to saturate at a constant value in the large-$\mathrm{bar}{n}$ limit, with the onset of this saturation determined by $n_{b}$ and $\kappa t$, see Fig. 5.
Figure 5: Noisy displacement sensitivity of pure Gaussian states. Left: Sensitivity calculated from Eq. (75) (solid lines) for an initially pure Gaussian state with $\overline{n}=1$ and different thermal baths with average photon number $n_{b}$ (colors), as a function of the loss $\kappa t$. The red dashed line corresponds to the sensitivity for the ideal unitary evolution. Right: sensitivity scaling at $\kappa t=0.1$.
@@ IV.2.1 Homodyne measurements
and analogously for $\max_{\varepsilon}(\chi^{-2})^{\min}_{\mathrm{hom,fix}}$ and $\max_{\varepsilon}(\chi^{-2})^{\mathrm{avg}}_{\mathrm{hom,fix}}$ with prefactors 0 and $1/2$, respectively. Since this maximum over $\varepsilon$ is achieved by choosing $\bm{w}$ along the direction of minimal phase-space variance, this means that, if $\phi$ is unknown, the above figures of merit are maximized by the measurement with the smallest variance. Similarly, we could easily obtain the minimum or average over $\varepsilon$. The evaluation of $(\chi^{-2})^{\max}_{\mathrm{hom,fix}}[\hat{\rho},\hat{r}(\varepsilon)]$ and $(\chi^{-2})^{\max}_{\mathrm{hom,max}}[\hat{\rho}$] for a selection of relevant families of states can be found in Table 4.
@@
Coherent $|\alpha\rangle$ 22 22
@@
[…]
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$, $\alpha\in\mathbb{C}$, $\left|\alpha\right|\gtrsim 2$ $\dfrac{2}{1+4\overline{n}^{2}\cos^{2}\varepsilon}$ 22
@@
Table 4: Displacement sensitivity attainable by linear quadrature measurements. Sensitivities computed according to Eqs. (81,84), corresponding to cases where the measurement direction $\varepsilon$ is chosen before the generator is known.
@@
$(\chi^{-2})^{\mathrm{avg}}_{\mathrm{hom}}[\hat{\rho}$] $=\frac{1}{2\pi}\int_{0}^{2\pi}\text{d}\phi\,\chi^{-2}_{\mathrm{hom}}[\hat{\rho},\hat{r}(\phi)]=\frac{\frac{1}{2}\left((\Delta\hat{x})^{2}_{\hat{\rho}}+(\Delta\hat{p})^{2}_{\hat{\rho}}\right)}{(\Delta\hat{x})^{2}_{\hat{\rho}}(\Delta\hat{p})^{2}_{\hat{\rho}}-\,\text{Cov}[\hat{x},\hat{p}]_{\hat{\rho}}^{2}}\;,$ (90)
@@
[…]
@@
Coherent $|\alpha\rangle$ 22 22 22
@@
Fock superposition
$(|0\rangle+e^{i\gamma}|n\rangle)/\sqrt{2}$
$\begin{cases}2\qquad&n=1\\ 2(3+\sqrt{2})/7&n=2\\ 2/(1+n)&\text{otherwise}\end{cases}$ $\begin{cases}\dfrac{2}{(3+\sqrt{2})}\qquad&n=2\\ \dfrac{2}{(1+n)}&\text{otherwise}\end{cases}$ $\begin{cases}3/2\qquad&n=1\\ 6/7&n=2\\ 2/(1+n)&\text{otherwise}\end{cases}$
Cat $\mathcal{N}(|\alpha\rangle+e^{i\gamma}|-\alpha\rangle)$
$\alpha\in\mathbb{C}$, $\left|\alpha\right|\gtrsim 2$
22 $\dfrac{2}{1+4\overline{n}}$ $\dfrac{2+4\overline{n}}{1+4\overline{n}}$
Compass
$\mathcal{N}(|\alpha\rangle+|-\alpha\rangle+|i\alpha\rangle+|-i\alpha\rangle)$
$\dfrac{2}{1+2\overline{n}}$ $\dfrac{2}{1+2\overline{n}}$ $\dfrac{2}{1+2\overline{n}}$
Table 5: Displacement sensitivity attainable by linear quadrature measurements. Sensitivities computed according to Eqs. (88,89,90), corresponding to cases where the measurement direction $\varepsilon$ is chosen after the generator is known.
@@ IV.2.2 Photon number measurements
where we used that $\hat{n},\hat{x}]=-i\hat{p}$ and $\hat{n},\hat{p}]=i\hat{x}$.
@@
$(\chi^{-2})^{\mathrm{avg}}_{\mathrm{num}}[\hat{\rho}]=\frac{1}{2\pi}\int_{0}^{2\pi}d\phi\chi^{-2}[\hat{\rho},\hat{r}(\phi),\hat{n}]=\frac{1}{2}\frac{\langle\hat{x}\rangle_{\hat{\rho}}^{2}+\langle\hat{p}\rangle_{\hat{\rho}}^{2}}{(\Delta\hat{n})^{2}_{\hat{\rho}}}\;.$ (94)
@@ IV.2.3 Measurement of higher-order moments
$\left\{\begin{aligned} \mathbf{\hat{r}}^{(1)}&=\mathbf{\hat{r}}=\{\hat{x},\hat{p}\}^{T},\\ \mathbf{\hat{r}}^{(2)}&=\{\hat{x}^{2},\frac{\hat{x}\hat{p}+\hat{p}\hat{x}}{2},\hat{p}^{2}\}^{T},\\ \mathbf{\hat{r}}^{(3)}&=\{\hat{x}^{3},\frac{\hat{x}^{2}\hat{p}+\hat{x}\hat{p}\hat{x}+\hat{p}\hat{x}^{2}}{3},\frac{\hat{x}\hat{p}^{2}+\hat{p}\hat{x}\hat{p}+\hat{p}^{2}\hat{x}}{3},\hat{p}^{3}\}^{T},\\$ $\vdots\end{aligned}\right$. (96)
@@
$\hat{M}_{\mathrm{opt}}=\cos(\phi)\left[-(2n+1)\hat{x}+\hat{x}^{3}+\frac{\hat{x}\hat{p}^{2}+\hat{p}\hat{x}\hat{p}+\hat{p}^{2}\hat{x}}{3}\right]-\sin(\phi)\left[-(2n+1)\hat{p}+\hat{p}^{3}+\frac{\hat{x}^{2}\hat{p}+\hat{x}\hat{p}\hat{x}+\hat{p}\hat{x}^{2}}{3}\right]\;.$ (97)
@@ IV.2.5 Parity operators and Wigner function measurements
according to Eq. (47), the moment-based sensitivity from its average value coincides with the classical Fisher information of the full probability for obtaining the results $1$ or 1-1 respectively,
@@
We note that there are points in phase space where Eq. (105) is indeterminate. The reason for this is because the Wigner function is a bounded function, and therefore both numerator and denominator of Eq. (105) are zero at its extremal points. In other words, these indeterminate values happen at points in phase space where a parity measurement gives $\pm 1$, such as for Fock states. However, we note that this indeterminate sensitivity is a mathematical artifact, that cannot be observed in experiments. In practice there will always be some finite amount of noise that prevents from observing $(\Delta\hat{M}(\beta))^{2}_{\hat{\rho}}=0$, and thus from having a zero denominator in Eq. (105). To describe this effect we add to the denominator of Eq. (105) a small positive number $\varepsilon$ representing measurement noise, namely we write
@@
Figure 6 shows the displacement sensitivity for parity measurements of Fock states, while Figures 7, 8 discuss superpositions of two Fock states. Figure 9 is for Gaussian states, while Figure 10 is for cats and compass states.
Figure 6: Displacement sensitivity of Fock states for parity measurements. Left panel: slices for $p=0$ of the Wigner functions for the first four Fock states. Right panel: sensitivity calculated from Eq. (105) (solid line), compared to the QFI bound $4n+2$ (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line).
Figure 7: Displacement sensitivity of superpositions of two Fock states for parity measurements. Top row: Wigner functions of different Fock state superpositions. Middle row: sensitivity to displacements obtained through Eq. (105). The displacement direction is optimized in order to maximise the sensitivity, and is indicated by coloured lines. The black dot indicate the point of maximum sensitivity (note that states with discrete rotation symmetry have several points with equal maximum sensitivity). Bottom row: sensitivity along the slices indicated by the lines in the two top rows calculated from Eq. (105) (solid line), compared to the QFI bound $4n+2$ (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line). Note that these two plots are not symmetric around zero, and that the QFI bounds are saturated for $n$ even or odd respectively, see Fig. 8.
Figure 8: Displacement sensitivity of superpositions of two Fock states for parity measurements. Sensitivity calculated from Eq. (105) (solid line), compared to the QFI bound $4n+2$ (dashed lines), and to the sensitivity achievable in the presence of noise with $\epsilon=10^{-2}$ (dotted line). Note that the QFI bounds are saturated for $n$ even or odd respectively.